OpenStudy (anonymous):

1/7(y+7)=2y with steps please

OpenStudy (mew55):

Distribute 1/7(y+7)=2y 1/7y+1=2y *1/7(7)=7/7=1 subtract 1/7 in both sides 1=13/7y Divide both sides by 13/7 y=7/13